# A Survey on the Applications of Zero-Knowledge Proofs

### v1

@@ IV-B3 Applications
zkSync Era [69] is a zkEVM rollup developed by Matter Labs. Identical to the previously mentioned methodology, zkSync Era utilizes ZK-SNARKs to provide validity proofs of off-chain computation. Smart contracts can be written in Solidity or Vyper and called using the same clients as other EVM-compatible chains, thus making zkSync EVM-compatible. The zksolc compiler used on Era generates bytecode with optimizations in order to make operations more amenable to proof generation, using LLVM as an intermediate representation before executing zkEVM assembly code. Consequently, there are multiple differences from Ethereum, with many EVM opcodes having modified implementation rules. zkSync Era is the first EVM-compatible chain to implement native account abstraction, which is a system of smart-contract-based accounts with arbitrary logic, first introduced in EIP-4337. Thus, every user account on zkSync Era can utilize smart accounts with their existing externally owned account (EOA).
@@
Matter labs also provides an open framework for deploying additional modular chains similar to zkSync Era called the ZK stack [69]. A modular chain from the ZK stack, called a hyperchain, runs a separate instance of the zkSync zkEVM and settles transactions on Ethereum’s Layer 1. Hyperchains are linked via Hyperbridges, ensuring asset transfer capabilities. While anyone can deploy Hyperchains, ensuring trust and full interoperability requires using the zkEVM engine from the ZK Stack, which powers zkSync Era. This uniformity in ZKP circuits across Hyperchains guarantees inherited security from Layer 1 without additional trust assumptions. Hyperchains implement a modular approach, allowing developers to choose or create their own blockchain components, except for the zkEVM core. Various options for sequencing transactions are available, ranging from centralized to decentralized sequencers, and even external protocols for customizing Hyperchain sequencing. Each Hyperchain can also manage its data availability (DA) policy, for example a “Validium” architecture which stores state data off-chain rather than posting the calldata to L1, providing flexibility tailored to specific needs.
Polygon zkEVM [70], built by Polygon Labs, aims to offer full EVM-equivalency, with no separate compiler. Thus, ZK proving circuits verify most EVM opcodes as they are, with a few carrying minor differences that do not impact the developer experience. Because of this inherent equivalency and support of EVM opcodes, developers can deploy their existing L1 smart contracts directly to the Polygon zkEVM rollup, with no necessary tweaks. Much like the zkSync stack, Polygon has the Chain Development Kit (CDK) [71], which allows developers to deploy application-specific chains as validiums using the Polygon zkEVM.
@@
Scroll [72] represents another approach within the zkEVM landscape, focusing on EVM compatibility with necessary adaptations for zero-knowledge proofs. Scroll’s zkEVM modifies certain EVM opcode behaviors to fit the ZK-proof framework while maintaining the ability for developers to write and deploy Solidity contracts, with no custom compiler. Although these modifications alter how some operations are handled compared to Ethereum, they are clearly documented, ensuring that developers can account for these changes during smart contract development. The tailored modifications aim to preserve the core experience of Ethereum smart contract interaction within the constraints of a ZK rollup environment.
@@
StarkNet [29], built by StarkWare, stands out as a unique ZK-rollup, fundamentally distinguished by its use of STARKs (Scalable Transparent ARguments of Knowledge), which offer quantum resistance and do not require a trusted setup, over SNARKs. Unlike other ZK-rollups, StarkNet utilizes Cairo – a specialized programming language – instead of Solidity. Cairo programs are compiled into Sierra, a safe intermediate representation, and subsequently into Cairo assembly (Casm) for execution by the StarkNet OS virtual machine. Use of the Cairo programming language and this two-step compilation process is necessary to bridge the gap between smart contract execution and the polynomial constraints of STARK proofs, which in turn validate StarkNet’s block execution. Like zkSync, StarkNet’s native account abstraction (AA) sets it apart by making all accounts ERC-4337 smart accounts with no externally owned accounts (EOAs) [74].
@@
StarkEx [75], also developed by StarkWare, is a specialized Layer 2 scalability service utilizing STARK proofs for high-throughput, low-latency applications on Ethereum. Unlike StarkNet, StarkEx is not a standalone blockchain, but a service specifically tailored for certain use cases like decentralized exchanges (DEXs) and NFT platforms. It allows these applications to define their own logic off-chain and post transactions to the service, which then generates STARK proofs attesting to the validity of transaction batches. These proofs are submitted and verified on L1. StarkEx also offers various data availability modes - ZK-Rollup, Validium, and Volition. This flexibility allows applications to optimize with any mix of on-chain and off-chain components.
@@
Aztec Network [76] is a Layer 2 rollup which focuses on privacy preservation through the Noir programming language. Noir, a Rust-based DSL for building ZK applications, simplifies the creation of ZKPs by abstracting the cryptographic process while retaining the robustness of circuit-building languages [77]. Smart contracts in Aztec, utilizing Noir, can have both public and private elements. These contracts are defined as sets of functions, both public and private, written as Noir circuits. Each function, representing a zk-SNARK verification key, operates on the contract’s public and private state. Noir’s compilation process is unique: it doesn’t compile directly to circuits but to an Abstract Circuit Intermediate Representation (ACIR), which can then be compiled into an arithmetic circuit or R1CS, depending on the proving system being targeted [39]. In Aztec, the sequencer aggregates transactions into a block, generates state update proofs, and posts them to the Ethereum rollup contract. This architecture, while similar to other Layer 2 networks, differs notably in its handling of private state. The private execution environment within Aztec safeguards sensitive operations and data, ensuring that private information remains confidential [76]. This architecture is outlined in Figure 10.
Figure 10: Aztec’s high-level network architecture, redrawn based on [76]
IV-C Blockchain Interoperability
IV-C1 Motivation and Definition
@@
With a highly fragmented landscape of different blockchain technologies, including Layer 1 chains and additional modular layers built atop them, there has arisen the need for seamless composability among blockchains. For example, a significant pain point in blockchain user experience is the struggle of trying to bridge tokenized assets from one blockchain to another. This problem extends past cross-chain transactions and financial asset liquidity, as it also concerns general message passing and data storage across fragmented blockchain networks. Even for Layer 2 rollups built atop the same Layer 1, each rollup’s state is a separate data moat, which results in the same fragmentation problem even in the case of shared transaction finality and consensus security. Cross-chain composability, also known as blockchain interoperability, has many innovative solutions, including those that utilize ZKP for succinct verifiable computation.
IV-C2 Methodology
@@
Zero-knowledge proofs are used to verify the occurrence of a state change or block execution on one chain to another chain. This property allows protocol developers to coordinate application logic across multiple blockchains, letting users instantly transact and pass data between different networks. Typically, there is some middleware that generates the validity proofs to be verified on the receiving blockchain, where a smart contract will verify the proof and execute corresponding application logic, according to the application’s specification. For example, asset bridges lock up tokens on one chain and mint them on another, allowing users to transfer liquidity between chains. Further use cases enable cross-chain DAO voting, Non-Fungible-Tokens (NFTs), and more.
IV-C3 Applications
@@
The zkBridge protocol [78], originally published as academic research and later implemented by Polyhedra[79], operates through a modular design that separates application-specific logic from the core functionality of relaying block headers. This core functionality is provided by a block header relay network, which is trusted only for liveness. This network relays block headers of one blockchain along with zk-SNARK correctness proofs to an updater contract on another blockchain. The updater contract is responsible for maintaining a list of recent block headers from the sender chain, verifying proofs submitted by relay nodes, and updating the list accordingly. On the receiver blockchain, the updater contract provides an application-agnostic API that enables application smart contracts to obtain the latest block headers of the sender blockchain. This enables them to build application-specific logic on top of this information. Applications utilizing zkBridge generally deploy a pair of contracts: a sender contract on blockchain 1 and a receiver contract on blockchain 2. The receiver contract can call the updater contract to access block headers of blockchain 1, which it can then use to execute application-specific tasks.
Figure 11: zkBridge architecture when used for a cross-chain token transfer, redrawn based on [78]
@@
IV-D Blockchain Storage
IV-D1 Motivation and Definition
@@
Blockchain storage is essentially a way to save data in a decentralized network using the properties of the blockchain. In order to protect the saved data, blockchain uses data structures such as Merkel trees and Merkel Patricia trees. What is special about these data structures is that specific proofs can be constructed to show that a particular piece of data is contained in the structure and that once a single piece of data is changed, the entire structure changes drastically as well. This essential property of blockchain ensures data integrity and invariance. However, as the quantity of data contained in a proof rises, so does the size of the proof. As a result, the cost of validating such proofs on the chain rises, rendering ordinary inclusion proofs economically unsustainable in most circumstances. A more scalable option is now preferred, whereby the blockchain does not have to store large amounts of data but only smaller references to data stored in off-chain platforms. Zero-knowledge proofs make this possible: data and computations can be stored off-chain, and ZK Proofs can be used to communicate a summary of these operations to the main chain concisely, efficiently, and without trust.
IV-D2 Methodology
@@
Zero-knowledge proofs could be used to minimize the cost of activities associated with verifying the inclusion of data in massive datasets and to validate that the procedure was completed correctly. The prover carries out the necessary calculation and generates proof that proves its correctness. The verifier’s job is to validate the proof’s validity without redoing the whole analysis. Because of this feature, the prover only requires access to a subset of the data, such as some nodes of a Merkle tree, rather than the entire dataset. This paradigm shift is crucial because it offers a practical solution to lower the costs of employing proofs of inclusion in smart contracts, particularly when massive datasets are involved.
IV-D3 Applications
@@
To prove the data is indeed stored in the blockchain, Herodotus [81] gives a new method called Storage Proofs to enable on-chain data access. It is essentially an on-chain accumulator that uses cryptographic means to improve access and verification of historical data on the Ether blockchain. The solution utilizes StarkWare’s STARK proofs [82]. This allows users to validate data from any point on the Ethernet blockchain without the need for a third party. It combines proofs of inclusion (for verifying the existence of data) and proofs of computation (for proving the execution of multi-step workflows). These proofs are essential for verifying the integrity and correctness of one or more components of a large data set, thus ensuring that the data has not only been stored, but also remains untampered with and accurate over time.
Filecoin [83] has deployed a considerable zk-SNARK network, which uses the unused hard drive space of users around the world to store files. Using zk-SNARKs to generate the proofs, the resulting proofs are small, and the verification process is swift (and thus, cheap). For example, proofs that typically would require hundreds of kilobytes to verify can instead be compressed to just 192 bytes with zk-SNARKs. This significant reduction in proof size not only accelerates the verification process but also reduces the computational and financial costs associated with it, making blockchain storage both more scalable and accessible.
IV-E Smart Contract/Transaction Privacy
IV-E1 Motivation and Definition
@@ IV-E2 Methodology
The core methodology in ZKP smart contract privacy centers around on-chain proof verification within a smart contract framework. Proof generation typically occurs off-chain due to its computationally intensive nature and the need for privacy in computation or the processing of sensitive data. Once a proof is generated, it is submitted to the blockchain. The contract assesses the proof against the protocol’s predefined rules, and then acts accordingly if deemed valid, which could involve updating the blockchain’s state, executing transactions, or any other protocol-specific actions.
IV-E3 Applications
Tornado Cash [42] employs zero-knowledge proofs for enabling transaction privacy and mixing on Ethereum. Its core mechanism revolves around depositing Ether into a smart contract and withdrawing it in a manner that severs the link between the source and destination. The protocol utilizes zk-SNARKs to prove the legitimacy of withdrawals without revealing the original deposit’s details. Users deposit Ether, generating a cryptographic commitment added to a Merkle tree within the contract. For withdrawal, a user generates a zk-SNARK proof that they own a leaf in this tree without revealing which one. This proof, once verified by the smart contract, allows the withdrawal of Ether to a new address, ensuring that the transaction’s privacy is maintained by obscuring the link between the deposit and withdrawal addresses. This method effectively creates a privacy layer, allowing users to transact anonymously within the public Ethereum blockchain.
@@ IV-E3 Applications
Due to its rampant use for illicit financial transactions such as money laundering, Tornado Cash has been heavily sanctioned and banned in countries like the US. This has given rise to the research problem of regulatory-compliant privacy solutions, which has been implemented in the Privacy Pools project, as delineated in [84]. Privacy Pools extends the usage of zk-SNARKs in other privacy solutions like Zcash and Tornado Cash. Rather than simply generating a ZKP to prove that a withdrawal attempt is linked to a specific deposit, users prove membership in a specific association set, which is a collection of transaction references, represented as a Merkle tree, from which a user’s funds could have originated. Users define their set by providing the Merkle root as a public input. When withdrawing funds, the user does not directly prove a link to a specific transaction. Instead, they utilize zk-SNARKs to generate proofs that validate their funds’ origin from within this predefined, public set of transactions. This approach allows users to demonstrate a connection to a pool of transactions without revealing the exact source, maintaining privacy while introducing an element of transparency.
The use of association sets in Privacy Pools addresses the challenge of aligning transaction privacy with regulatory compliance. By proving membership in a specific association set, users demonstrate that their funds originate from a group of transactions that are not flagged as high-risk or associated with illicit activities.
@@
Penumbra [85] is a fully private proof-of-stake network and decentralized exchange within the Cosmos ecosystem, offering a unique approach to transaction privacy and cross-blockchain interoperability using zero-knowledge proofs. Penumbra connects to the Cosmos blockchain ecosystem through IBC (the inter-blockchain communication protocol) and maintains all value in a multi-asset shielded pool, inspired by the Zcash Sapling design. This allows for private transactions in any IBC-compatible asset. All transactions on Penumbra are private by default, enabled by zk-SNARKS which validate the correctness and legitimacy of transactions but shield the sender, receiver, and amount. Penumbra’s decentralized exchange, called ZSwap, supports sealed-bid batch auctions and concentrated liquidity similar to Uniswap v3. This architecture prevents frontrunning and ensures that only the net flow across a pair of assets is revealed in each block. For cryptographic primitives, Penumbra utilizes BLS12-377 as the proving curve, which is compatible with the Groth16 proving system used. Penumbra may change to PLONK in the future.
A novel privacy problem exists within the space of on-chain decentralized autonomous organizations (DAOs), which are blockchain-based governance systems powered by smart contracts. Because of the transparent nature of public blockchains, DAO treasuries are completely public, which is a big issue for certain types of auctions. In a blog and research project by Griffin Dunaif and Dan Boneh, the authors design a system for a private DAO protocol utilizing ZKP [86]. The protocol utilizes a master contract deployed on the Ethereum network to manage multiple DAOs. This contract allows anyone to send funds to a DAO, but only the DAO manager can withdraw them.
@@
The life cycle of a DAO in this system comprises three steps: creation, deposit, and withdrawal. During DAO creation, the manager establishes the DAO without any on-chain transactions, and posts a Schnorr public key on the DAO website. To contribute funds to the DAO, users compute the Merkle tree leaf using the DAO public key, in which the deposit is recorded in the master contract. The DAO manager is able to privately monitor deposits and keep track of the DAO treasury using their secret key. When withdrawing funds, the DAO manager again uses this secret key within a SNARK proof to show that a specific batch of deposit leaves in the Merkle tree belong to the DAO, without publicly revealing the secret key. After verifying the proof, the master contract releases the withdrawal amount to the DAO manager.
@@
IV-F Blockchain-Based Proof of Identity
IV-F1 Motivation and Definition
Because of the aforementioned qualities of ZKP, there exists the unique use-case of a user proving group membership or identity without revealing any sensitive information about their identity, thus marrying the imperatives of authentication and privacy. While this can be implemented outside of the blockchain context (as will be covered in the “Proof of Identity” section below), there exist many identity-proving applications using blockchain smart contracts and execution layers. This application domain can be defined as Blockchain-Based Proof of Identity, which takes advantage of privacy-preserving ZKP.
IV-F2 Methodology
The core methodology behind blockchain-based proof of identity systems is the use of zero-knowledge proofs to enable users to prove membership in certain identity groups or ownership of credentials without revealing the actual identity information. This is achieved through cryptographic commitments to a user’s information while keeping it secret. A privacy-preserving ZK identity proof can be easily verified by an on-chain smart contract, and this can act as a private credential system secured by the blockchain and open doors to a multitude of decentralized applications, without sensitive information compromise.
IV-F3 Applications

### v2

@@ 4.2.3 Applications
Table 8 summarizes representative design points to aid interpretation of the following systems: zkSync Era [58] is a zkEVM rollup developed by Matter Labs. Identical to the previously mentioned methodology, zkSync Era utilizes zkSNARKs to provide validity proofs of off-chain computation. Smart contracts can be written in Solidity or Vyper and called using the same clients as other EVM-compatible chains, thus making zkSync EVM-compatible. The zksolc compiler used on Era generates bytecode with optimizations in order to make operations more amenable to proof generation, using LLVM as an intermediate representation before executing zkEVM assembly code. Consequently, there are multiple differences from Ethereum, with many EVM opcodes having modified implementation rules. zkSync Era is an early EVM-compatible chain to implement native account abstraction, a system of smart-contract-based accounts with arbitrary logic first introduced in EIP-4337. Thus, every user account on zkSync Era can use smart accounts with their existing externally owned accounts (EOAs).
@@
Matter labs also provides an open framework for deploying additional modular chains similar to zkSync Era called the ZK stack [58]. A modular chain from the ZK stack, called a hyperchain, runs a separate instance of the zkSync zkEVM and settles transactions on Ethereum’s Layer 1. Hyperchains are linked via Hyperbridges, enabling asset transfers. While anyone can deploy Hyperchains, ensuring trust and full interoperability requires using the zkEVM engine from the ZK Stack, which powers zkSync Era. This uniformity in ZKP circuits across Hyperchains guarantees inherited security from Layer 1 without additional trust assumptions. Hyperchains implement a modular approach, allowing developers to choose or create their own blockchain components, except for the zkEVM core. Various options for sequencing transactions are available, ranging from centralized to decentralized sequencers to external protocols for customizing Hyperchain sequencing. Each Hyperchain can also manage its data availability (DA) policy, for example, a “Validium” architecture which stores state data off-chain rather than posting the calldata to L1, providing flexibility tailored to specific needs.
@@
Polygon zkEVM [78], built by Polygon Labs, aims to offer full EVM-equivalency, with no separate compiler. Thus, ZK proving circuits verify most EVM opcodes as they are, with a few that carry minor differences that do not affect the developer experience. Because of this inherent equivalence and support for EVM opcodes, developers can deploy their existing L1 smart contracts directly to the Polygon zkEVM rollup with no necessary tweaks. Much like the zkSync stack, Polygon has the Chain Development Kit (CDK) [77], which allows developers to deploy application-specific chains as validiums using the Polygon zkEVM. Scroll [88] represents another approach within the zkEVM landscape, focusing on EVM compatibility with necessary adaptations for zero-knowledge proofs. Scroll’s zkEVM modifies certain EVM opcodes to fit the ZK-proof framework while maintaining the ability for developers to write and deploy Solidity contracts with no custom compiler. Although these modifications alter how some operations are handled compared to Ethereum, they are clearly documented, allowing developers to account for them during smart contract development. The tailored modifications aim to preserve the core experience of Ethereum smart contract interaction within the constraints of a ZK rollup environment.
@@
StarkNet [92], built by StarkWare, differs from other ZK-rollups in that it uses STARK proofs rather than SNARK-based schemes. StarkNet uses Cairo, a specialized programming language, rather than Solidity. Cairo programs are compiled into Sierra, a safe intermediate representation, and subsequently into Cairo assembly (Casm) for execution by the StarkNet OS virtual machine. This two-step compilation process translates smart contract execution into the polynomial constraints required by STARK proofs, which validate StarkNet’s block execution. Like zkSync, StarkNet adopts a native account-abstraction model in which accounts are smart-contract-based rather than traditional EOAs [95].
@@
StarkEx [96], also developed by StarkWare, is a specialized Layer 2 scalability service utilizing STARK proofs for high-throughput, low-latency applications on Ethereum. Unlike StarkNet, StarkEx is not a standalone blockchain but a service tailored for specific use cases, such as decentralized exchanges (DEXs) and NFT platforms. It allows these applications to define their own logic off-chain and post transactions to the service, which then generates STARK proofs attesting to the validity of transaction batches. These proofs are submitted and verified on L1. StarkEx also offers various data availability modes: ZK-Rollup, Validium, and Volition. This flexibility allows applications to optimize with any mix of on-chain and off-chain components.
@@
Aztec Network [9] is a Layer 2 rollup which focuses on privacy preservation through the Noir programming language [70]. Smart contracts in Aztec using Noir can include both public and private elements. These contracts are defined as sets of functions, both public and private, written as Noir circuits. Each function, representing a zk-SNARK verification key, operates on the contract’s public and private state. In Aztec, the sequencer aggregates transactions into a block, generates state update proofs, and posts them to the Ethereum rollup contract. This architecture, while similar to other Layer 2 networks, differs notably in how it handles private state. The private execution environment within Aztec safeguards sensitive operations and data, ensuring that private information remains confidential [9].
4.3 Blockchain Interoperability
4.3.1 Motivation and Definition
@@
With a highly fragmented landscape of different blockchain technologies, including Layer 1 chains and additional modular layers built atop them, there has arisen the need for composability among blockchains. For example, a significant pain point in the blockchain user experience is the difficulty of bridging tokenized assets from one blockchain to another. This problem extends past cross-chain transactions and financial asset liquidity, as it also concerns general message passing and data storage across fragmented blockchain networks. Even for Layer 2 rollups built atop the same Layer 1, each rollup’s state constitutes a separate data moat, resulting in the same fragmentation problem even with shared transaction finality and consensus security. Cross-chain composability, also known as blockchain interoperability, has produced several approaches, including those that leverage ZKPs for succinct verifiable computation.
4.3.2 Methodology
@@
Zero-knowledge proofs are used to verify the occurrence of a state change or block execution on one chain to another chain. This property enables protocol developers to coordinate application logic across multiple blockchains, allowing users to instantly transact and pass data between networks. Typically, middleware generates validity proofs that are verified on the receiving blockchain, where a smart contract verifies the proof and executes the corresponding application logic according to the application’s specification. For example, asset bridges lock up tokens on one chain and mint them on another, allowing users to transfer liquidity between chains. Further use cases include cross-chain DAO voting and Non-Fungible Tokens (NFTs).
4.3.3 Applications
@@
The zkBridge protocol [105] illustrated in Figure 7, operates through a modular design that separates application-specific logic from the core functionality of relaying block headers. This core functionality is provided by a block header relay network, which is trusted only for liveness. This network relays block headers from one blockchain, along with zkSNARK correctness proofs, to an updater contract on another blockchain. The updater contract is responsible for maintaining a list of recent block headers from the sender chain, verifying proofs submitted by relay nodes, and updating the list accordingly. On the receiver blockchain, the updater contract provides an application-agnostic API that enables application smart contracts to obtain the latest block headers of the sender blockchain. This enables them to build application-specific logic on top of this information. Applications utilizing zkBridge generally deploy a pair of contracts: a sender contract on blockchain 1 and a receiver contract on blockchain 2. The receiver contract can call the updater contract to access block headers of blockchain 1, which it can then use to execute application-specific tasks.
Figure 7: zkBridge architecture for a cross-chain token transfer
@@
4.4 Blockchain Storage
4.4.1 Motivation and Definition
@@
Blockchain storage is essentially a way to save data in a decentralized network using the properties of the blockchain. In order to protect the saved data, blockchain uses data structures such as Merkle trees and Merkle Patricia trees. What is special about these data structures is that specific proofs can be constructed to show that a particular piece of data is contained in the structure and that, once a single piece of data is changed, the entire structure changes drastically. This essential property of blockchain ensures data integrity and invariance. However, as the quantity of data contained in a proof rises, so does the size of the proof. As a result, the cost of validating such proofs on the chain rises, rendering ordinary inclusion proofs economically unsustainable in most circumstances. A more scalable option is now preferred, in which the blockchain does not have to store large amounts of data but instead only small references to data stored on off-chain platforms. Zero-knowledge proofs make this possible: data and computations can be stored off-chain, and ZK Proofs can be used to succinctly and efficiently communicate a summary of these operations to the main chain without trust.
4.4.2 Methodology
@@
Zero-knowledge proofs could be used to minimize the cost of activities associated with verifying the inclusion of data in massive datasets and to validate that the procedure was completed correctly. The prover carries out the necessary calculation and generates a proof of its correctness. The verifier’s job is to verify the proof without redoing the entire analysis. Because of this feature, the prover only requires access to a subset of the data, such as a subset of nodes in a Merkle tree, rather than the entire dataset. This shift is significant because it offers a practical way to lower the costs of employing proofs of inclusion in smart contracts, particularly when massive datasets are involved.
4.4.3 Applications
@@
To prove the data is indeed stored in the blockchain, Herodotus [43] gives a new method called Storage Proofs to enable on-chain data access. It is essentially an on-chain accumulator that uses cryptographic means to improve access and verification of historical data on the Ethereum blockchain. The solution utilizes StarkWare’s STARK proofs [11]. This allows users to validate data from any point on the Ethereum blockchain without a third party. It combines proofs of inclusion for verifying the existence of data and proofs of computation for proving the execution of multi-step workflows. These proofs are essential for verifying the integrity and correctness of one or more components of a large data set, thus ensuring that the data has not only been stored but also remains untampered with and accurate over time.
Filecoin [82] has deployed a large-scale zkSNARK network that uses unused hard drive space from users worldwide to store files. Using zkSNARKs to generate proofs results in small proofs, and the verification process is swift and thus cheap. In some deployed constructions, zkSNARKs can reduce proof artifacts from much larger alternatives to very small constant-size proofs, substantially lowering verification overhead and making blockchain storage both more scalable and accessible.
4.5 Smart Contract/Transaction Privacy
4.5.1 Motivation and Definition
@@ 4.5.2 Methodology
The core methodology in ZKP smart contract privacy centers around on-chain proof verification within a smart contract framework. Proof generation typically occurs off-chain due to its computationally intensive nature and the need for privacy in computation or the processing of sensitive data. Once a proof is generated, it is submitted to the blockchain. The contract assesses the proof against the protocol’s predefined rules, and then acts accordingly if deemed valid, which could involve updating the blockchain’s state, executing transactions, or any other protocol-specific actions. In practice, proof-system choice is driven by on-chain verification cost: Groth16 is widely used in deployed privacy contracts in part because its constant-size proofs can yield relatively low on-chain verification cost compared with larger-proof alternatives, whereas STARK proofs, though transparent and post-quantum, are typically larger and therefore often less attractive in gas-sensitive settings. Universal-setup schemes such as PLONK offer a middle ground–updatable structured reference strings and competitive proof sizes–and are increasingly adopted by projects that need circuit flexibility without per-application trusted setups.
4.5.3 Applications
Tornado Cash [75] employs zero-knowledge proofs for enabling transaction privacy and mixing on Ethereum. Statement proved: set membership in a Merkle tree of deposit commitments, combined with a nullifier check that prevents double-spending. Privacy goal: unlinking deposit and withdrawal addresses so that an observer cannot determine which deposit funds a given withdrawal. Engineering bottleneck: in Tornado Cash’s design, the effective anonymity set grows only as fast as new deposits arrive, and in typical implementations, on-chain verification for a withdrawal proof is on the order of a few hundred thousand gas. Limitations: timing correlation between deposits and withdrawals can degrade anonymity in practice, and the fixed-denomination design (e.g., 0.1, 1, 10, 100 ETH) introduces a UX constraint that limits flexible-amount transfers.
@@ 4.5.3 Applications
Tornado Cash’s deployment demonstrated that protocol-level unlinkability can attract both legitimate privacy demand and illicit use, leading to regulatory sanctions in several jurisdictions. This has given rise to the research problem of regulatory-compliant privacy solutions, which has been implemented in the Privacy Pools project, as delineated in [18]. Privacy Pools extends the usage of zkSNARKS in other privacy solutions like Zcash and Tornado Cash. Rather than simply generating a ZKP to prove that a withdrawal attempt is linked to a specific deposit, users prove membership in a specific association set, which is a collection of transaction references, represented as a Merkle tree, from which a user’s funds could have originated. Users define their set by providing the Merkle root as a public input. When withdrawing funds, the user does not directly prove a link to a specific transaction. Instead, they utilize zkSNARKS to generate proofs that validate their funds’ origin from within this predefined, public set of transactions. By proving membership in a curated association set, users can demonstrate that their funds do not originate from flagged addresses, aligning transaction privacy with compliance requirements without sacrificing the cryptographic unlinkability of individual transfers.
More broadly, the tension between privacy and accountability has spurred several complementary design patterns. Selective disclosure mechanisms allow authorized auditors to decrypt transaction metadata under predefined legal conditions while preserving unlinkability for ordinary observers [104]. Regulated set membership proofs extend the association-set idea by requiring withdrawals to satisfy compliance predicates–e.g., that funds have not passed through sanctioned addresses within a configurable lookback window [18]. Accountability-focused designs incorporate revocation or tracing capabilities that can be activated only through multi-party authorization, preventing unilateral surveillance while enabling lawful investigation [33]. These approaches remain at varying levels of maturity, and their interaction with jurisdiction-specific regulations is an active area of research.
@@
Penumbra [74] is a fully private proof-of-stake network and decentralized exchange within the Cosmos ecosystem, offering a unique approach to transaction privacy and cross-blockchain interoperability using zero-knowledge proofs. Penumbra connects to the Cosmos blockchain ecosystem through IBC (the inter-blockchain communication protocol) and maintains all value in a multi-asset shielded pool, inspired by the Zcash Sapling design. This allows for private transactions in any IBC-compatible asset. All transactions on Penumbra are private by default, enabled by zkSNARKS which validate the correctness and legitimacy of transactions but shield the sender, receiver, and amount. Penumbra’s decentralized exchange, called ZSwap, supports sealed-bid batch auctions and concentrated liquidity similar to Uniswap v3. This architecture prevents frontrunning and ensures that only the net flow across a pair of assets is revealed in each block. For cryptographic primitives, Penumbra utilizes BLS12-377 as the proving curve, which is compatible with the Groth16 proving system used. Penumbra may change to PLONK in the future.
@@
A novel privacy problem exists within the space of on-chain decentralized autonomous organizations (DAOs), which are blockchain-based governance systems powered by smart contracts. Because of the transparent nature of public blockchains, DAO treasuries are completely public, which is a big issue for certain types of auctions. In a blog and research project by Griffin Dunaif and Dan Boneh, the authors design a system for a private DAO protocol utilizing ZKP [24]. The protocol utilizes a master contract deployed on the Ethereum network to manage multiple DAOs. This contract allows anyone to send funds to a DAO, but only the DAO manager can withdraw them. The life cycle of a DAO in this system comprises three steps: creation, deposit, and withdrawal. During DAO creation, the manager establishes the DAO without any on-chain transactions, and posts a Schnorr public key on the DAO website. To contribute funds to the DAO, users compute the Merkle tree leaf using the DAO public key, in which the deposit is recorded in the master contract. The DAO manager is able to privately monitor deposits and keep track of the DAO treasury using their secret key. When withdrawing funds, the DAO manager again uses this secret key within a SNARK proof to show that a specific batch of deposit leaves in the Merkle tree belong to the DAO, without publicly revealing the secret key. After verifying the proof, the master contract releases the withdrawal amount to the DAO manager.
@@
4.6 Blockchain-Based Proof of Identity
4.6.1 Motivation and Definition
Because of the aforementioned qualities of ZKP, there exists the unique use-case of a user proving group membership or identity without revealing any sensitive information about their identity, thus marrying the imperatives of authentication and privacy. While this can be implemented outside of the blockchain context (as will be covered in the “Proof of Identity" section below), there exist many identity-proving applications using blockchain smart contracts and execution layers. This application domain can be defined as Blockchain-Based Proof of Identity, which takes advantage of privacy-preserving ZKP.
4.6.2 Methodology
The core methodology behind blockchain-based proof of identity systems is the use of zero-knowledge proofs to enable users to prove membership in certain identity groups or ownership of credentials without revealing the actual identity information. This is achieved through cryptographic commitments to a user’s information while keeping it secret. A privacy-preserving ZK identity proof can be easily verified by an on-chain smart contract, and this can serve as a private credential system secured by the blockchain, enabling decentralized applications without exposing sensitive information.
4.6.3 Applications